OpenStudy (anonymous):

pls help.simplify 1.gamma (a+b) 2.gamma (a+1) 3.gamma (a+b+1) will appreciate ur hlp

OpenStudy (anonymous):

some property of the gamma function? only thing i remember about it is that it is a generalization of factorial. so \(\Gamma(n)=(n-1)!\) for integer n that would make \(\Gamma(a+b+1)=(a+b)\Gamma(a+b)\) but other than that i am lost